Bearclaws72:

The main tank has a radius of 70 feet. What is the volume of the quarter-sphere sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit.

Bearclaws72:

Yee yee I don't know how to do this

misty34:

1) V=4/3pi*r^3 Is the volume formula 2) Plug in the numbers into the formula 3) You would get V= 4/3 * 3.14 * (70 )^ 3 = 1436027 4) 1/4 sphere= 1436027/4 5) You would get an answer of 359007 ft. ^3